An electrical company takes 120 hours with a standard deviation of 25 hours to wire a build complex. What is probability that the same company will take at least 115 hours to do the same job?
Answer by stanbon(75887)

An electrical company takes 120 hours with a standard deviation of 25 hours to wire a build complex. What is probability that the same company will take at least 115 hours to do the same job?
--------
z(115) = (115-120)/25 = -1/5
------
P(x >= 115) = P(z >= -1/5) = 0.5793
=============
Cheers,
Stan H.
=============